# Parameter-Efficient Continual Fine-Tuning: A Survey

### v1

@@ 7 Conclusion
In this work, we tackled a key limitation of current foundation models: their restricted ability to adapt to specific downstream tasks over time. To address this, we focused on two research fields aimed at enhancing adaptation in dynamic scenarios: Continual Learning and Parameter-Efficient Fine-Tuning. One one hand, CL enables models to incrementally learn from a continuous stream of tasks, while PEFT facilitates rapid adaptation to individual tasks. We explored the intersection of these two fields — which we called Parameter-Efficient Continual Fine-Tuning (PECFT) — and examined how their techniques can be combined to improve the efficiency and sustainability of large pre-trained models in evolving environments. We believe that bridging CL and PEFT has the potential to drive significant advancements in AI research, shaping the future of large-scale models. Through this review, we aim to provide researchers with an overview of existing methods in this emerging area, fostering the development of novel and effective solutions.
@@
During the preparation of this work the author(s) used ChatGPT in order to enhance readability and check grammar. After using this tool, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.
@@ Acknowledgements
Research partly funded by PNRR - M4C2 - Investimento 1.3, Partenariato Esteso PE00000013 - ”FAIR - Future Artificial Intelligence Research” - Spoke 1 ”Human-centered AI”, funded by the European Commission under the NextGeneration EU programme and Leonardo Labs.

### v2

@@ 7 Conclusion
In this work, we tackled a key limitation of current foundation models: their restricted ability to adapt to specific downstream tasks over time. To address this, we focused on two research fields aimed at enhancing adaptation in dynamic scenarios: Continual Learning and Parameter-Efficient Fine-Tuning. One one hand, CL enables models to incrementally learn from a continuous stream of tasks, while PEFT facilitates rapid adaptation to individual tasks. We explored the intersection of these two fields , which we called Parameter-Efficient Continual Fine-Tuning (PECFT), and examined how their techniques can be combined to improve the efficiency and sustainability of large pre-trained models in evolving environments. We believe that bridging CL and PEFT has the potential to drive significant advancements in AI research, shaping the future of large-scale models. Through this review, we aim to provide researchers with an overview of existing methods in this emerging area, fostering the development of novel and effective solutions.
@@
During the preparation of this work the authors used ChatGPT in order to enhance readability and check grammar. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.
After using this tool, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.
@@ Acknowledgements
Research partly funded by PNRR - M4C2 - Investimento 1.3, Partenariato Esteso PE00000013 - “FAIR - Future Artificial Intelligence Research” - Spoke 1 “Human-centered AI”, funded by the European Commission under the NextGeneration EU programme and Leonardo Labs.